# A Framework of SO(3)-equivariant Non-linear Representation Learning and its Application to Electronic-Structure Hamiltonian Prediction

### v3

@@ 4 Results and Analyses
We present experimental results of ablation study in Supplementary Information F. These fine-grained experiments assess the individual and combined contributions of the two core mechanisms of the proposed TraceGrad method on the six databases from the DeepH series. These mechanisms are the SO(3)-invariant trace supervision mechanism (Trace) at the label level and the gradient-based induction mechanism (Grad) at the representation layer. Results indicate that each mechanism can contribute individually to the performance. Moreover, their combination provides even better performance. This is because, on one hand, with the gradient-based induction mechanism as a bridge, the SO(3)-invariant features learned from the trace label can be transformed into the SO(3)-equivariant representations during inference; on the other hand, with trace label, the SO(3)-invariant network branch has a strong supervisory signal, enabling it to learn the intrinsic symmetry and complexity of the regression targets, enhancing the quality of SO(3)-invariant features and ultimately benefits the encoding of SO(3)-equivariant features. The value of such complementarity has been fully demonstrated in the experimental results.
5 Future Work
First, from a theoretical perspective, we can generalize SO(3) group symmetry to a broader range and more complex groups. This will enable deep learning frameworks to accommodate richer mathematical structures.
Second, from a methodological perspective, we plan to combine our framework with advanced Transformer [23] and KAN [24] architectures to continuously enhance the expressive capabilities of equivariant features.
Third, from the application perspective, the current methodology extends beyond predicting electronic-structure Hamiltonians and is capable of generally predicting multiple physical quantities and properties that require equivariance. Furthermore, the methodology is not confined to physical research alone and may find applications in areas such as robotics, autonomous vehicles, and motion tracking systems. For instance, in the vision tasks of autonomous vehicles [25], where there is a high demand for robustness against coordinate transformations of cameras, the mainstream approach has been to achieve approximate SO(3)-equivariance through data augmentation. However, this does not guarantee absolute reliability and safety. Our work holds great potential to construct deep models with high generalization performance under the premise of strict SO(3)-equivariance, potentially revolutionizing techniques in such tasks. Looking forward to future endeavors, the foundational and pivotal role of this work is clear.
6 Method
6.1 Encoding and Decoding Framework
As shown in Fig. 2, our encoding framework corresponds to a total of $K$ encoding modules, which are sequentially connected to form a cascading stack. For the $k$-th module ($1\leq k\leq K$), it first introduces an SO(3)-equivariant base encoder, e.g., the encoder from DeepH-E3 [8] or QHNet [7] which is composed of recently developed equivariant operators [26, 27] like linear scaling, direct products, direct sums, Clebsch-Gordan decomposition, gated activation, equivariant normalization, and etc., to encode the physical system’s initial representations, such as spherical harmonics [28], or representations passed from the previous neural layers, as equivariant feature $\textbf{f}^{(k)}$ in the current hidden layer. Next, we construct the feature $\textbf{v}^{(k)}=g_{\text{nonlin}}(\textbf{f}^{(k)})$, to achieve sufficient non-linear expressiveness while maintaining SO(3)-equivariance, where $g_{\text{nonlin}}(\cdot)$ is the function defined in Eq. (5). $s_{nonlin}(\cdot)$, i.e., $g_{\text{nonlin}}(\cdot)$’s built-in non-linear mechanism, which can be any differentiable non-linear neural module, is implemented here as a feed-forward module with a hidden layer size of $1024$, using Layer Normalization for stabilization and regularization, and employing SiLU as the non-linear mapping. To facilitate learning and optimization, we connect the input and output of $g_{\text{nonlin}}(\cdot)$ in a residual manner [29], serving as the output of the current encoding module, i.e., $\textbf{o}^{(k)}=\textbf{f}^{(k)}+\textbf{v}^{(k)}$. $\textbf{o}^{(k)}$ is then fed into the next encoder module to yield $\textbf{f}^{(k+1)}$, $\textbf{o}^{(k+1)}$, …., until the last encoder module, which is followed by the decoder modules.
For an expressive feature encoder handling a complex physical system, $\textbf{f}^{(k)}$ is usually not a feature of single degree but a direct-sum concatenation of series of components $\textbf{f}^{(k)l_{1}}$, $\textbf{f}^{(k)l_{2}}$, …, $\textbf{f}^{(k)l_{k}}$ at multiple degrees, i.e., $l_{1}$, $l_{2}$, …, $l_{k}$, where some components share the same degree while others differ. In this case, it becomes necessary to extend the decomposition operator, i.e., $\text{CGDecomp}(\cdot)$ in Eq. (5), to accommodate various components of degrees. Moreover, in the context of neural networks, introducing learnable parameters $W$ may improve the model capacity. Based on these considerations, when constructing the encoding module, the $\text{CGDecomp}(\cdot)$ operation in Eq. (5) can be expanded as $\text{CGDecomp}_{ext}(\cdot)$:
$u^{(k)}=\text{CGDecomp}_{\text{ext}}(\textbf{f}^{(k)}\otimes\textbf{f}^{(k)},0,W)=\sum_{1\leq i,j\leq k,l_{i}=l_{j}}W_{ij}\cdot\text{CGDecomp}(\textbf{f}^{(k)l_{i}}\otimes\textbf{f}^{(k)l_{j}},0)$ (6)
where $W$ represents learnable parameters. This expansion does not change the equivariance or invariance of the features. It is obvious that, the output ($u^{(k)}$) of $\text{CGDecomp}_{\text{ext}}(\cdot)$ is also SO(3)-invariant, just like the original $\text{CGDecomp}(\cdot)$. Thus, even with such an expansion, the feature $z^{(k)}=s_{\text{nonlin}}(u^{(k)})$ remains SO(3)-invariant and $\textbf{v}^{(k)}=\frac{\partial z^{(k)}}{\partial\mathbf{f}^{(k)}}$ remain SO(3)-equivariant.
[…]
Since complex non-linear transformations have already been incorporated in the encoding framework, the decoders can be simple: the SO(3)-equivariant decoder is implemented using an SO(3)-equivariant linear mapping [27], and the SO(3)-invariant decoder is implemented as a shallow feed-forward module.
6.2 Training and Inference
The training loss function is shown as the following:
$\begin{split}\min_{\theta}loss&=loss_{Q}+\mu(loss_{Q},loss_{T})\cdot loss_{T},\\ loss_{Q}&=MSE(\widehat{\mathbf{Q}},\mathbf{Q}^{*}),\\ loss_{T}&=MSE(\widehat{\mathbf{T}},\mathbf{T}^{*})\end{split}$ (7)
where $\theta$ denotes all of the learnable parameters of our framework, $\widehat{\mathbf{Q}}$, $\widehat{\mathbf{T}}$ and $\mathbf{Q}^{*}$, $\mathbf{T}^{*}$ respectively denote the predictions and labels of $\mathbf{Q}$ and $\mathbf{T}$. In order that the both of the SO(3)-equivariant and SO(3)-invariant branches can be stably converged, we introduce $\mu(loss_{Q},loss_{T})$, i.e., a function returning a coefficient to regularize the relative scale of the two loss terms:
where $No\_Grad(\cdot)$ denotes gradient discarding when calculating such coefficient, as this coefficient is only used for adjusting the relative scale of the two loss terms and should not itself be a source of training gradients, otherwise it would counteract $loss_{T}$ in Eq. (7). The introduction of $\mu(loss_{Q},loss_{T})$ effectively prevents the numerical disparity between the two loss terms to stabilize the training for both of them. All of the encoding and decoding modules are trained jointly by Eq. (7), and the training configurations are included in Supplementary Information D.
In our framework, the SO(3)-invariant non-linear features $z^{(k)}(1\leq k\leq K)$ learned with the help of $\mathbf{T}$ contribute to the learning of SO(3)-equivariant representations $\textbf{v}^{(k)}(1\leq k\leq K)$ and the prediction of $\mathbf{Q}$. To achieve this goal, in the training phase, both the SO(3)-invariant feature encoder and the SO(3)-invariant decoder, used for predicting $\mathbf{T}$ from $z^{(k)}$, need to be trained. While in the inference phase, since $\mathbf{T}$ is not the final regression target, the SO(3)-invariant decoder can be removed, yet the SO(3)-invariant encoder still should be retained, used to induce SO(3)-equivariant features.
6.3 Time Complexity
Typically, one of the key advantage of deep learning algorithms for electronic-structure calculation lies in their low computational complexity which can asymptotically scale as $O(N)$, where $N$ is the number of atoms of a system. This is a significant efficiency advantage compared to the $O(N^{3})$ computational complexity of traditional DFT method, enabling efficient and low-cost prediction of electronic structures for large-scale atomic systems. Our method improves the accuracy of Hamiltonian predictions within the deep learning paradigm while maintaining this advantage of asymptotic complexity. The baseline models we select, whether DeepH-E3 or QHNet, are efficient models with an asymptotic computational complexity of $O(N)$ as $N$ increases. The neural module $g_{nonlin}(\cdot)$ we introduce efficiently operates on each individual feature f from the baseline models to obtain the corresponding non-linear SO(3)-equivariant feature v, rather than relying on complex interactions between multiple features to derive new ones. This ensures that the computational complexity remains asymptotically $O(N)$, with good potential to efficiently handle large atomic systems.
7 Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant Number 12134012) and the Strategic Priority Research Program of Chinese Academy of Sciences (Grant Number XDB0500201).
Appendix A The Equivariance-Expressiveness Dilemma in Existing Deep Learning Methods
In this section, we introduce existing efforts to harmonize the deep learning paradigm with SO(3)-equivariance, analyzing the dilemma between strict SO(3)-equivariance and non-linear expressive capabilities. This dilemma leads to an unsatisfactory trade-off between strict equivariance and non-linear expressiveness, ultimately limiting the achievable performance and scopes of application for these regression models.
A.1 Approaches with Strict SO(3)-equivariance but Weak Expressiveness
Most existing methods have focused on achieving SO(3)-equivariance at the neural mechanism level by applying group theory-guaranteed SO(3)-equivariant operators [26, 27]. These operators, including linear scaling, element-wise sum, direct products, direct sums, and Clebsch-Gordan decomposition, have been used to build graph neural network architectures for tasks in 3D point cloud analysis [30], molecular dynamic simulation and property prediction [13, 12], as well as Hamiltonian prediction [31, 4, 8, 14]. However, these operators could not provide sufficient expressiveness to fit complex non-linear mappings. The majority of these operators are linear operations, and the only one (direct product) that introduces some non-linearity can only approximate a simple squaring function, and cannot be extended to more complex non-linear forms due to the huge computational burden and difficulty in convergence. Even worse, directly applying highly-expressive non-linear activation functions, such as $Sigmoid$, $Softmax$, and $SiLU$, to SO(3)-equivariant features may result in the loss of equivariance. To alleviate this issue, these methods introduced a gated activation mechanism that processes SO(3)-invariant features ($l=0$) using non-linear activations and employs them as linear gating coefficients for SO(3)-equivariant features ($l\geq 1$), aiming to enhance their expressive power while maintaining strict equivariance. However, from the perspective of SO(3)-equivariant features, this mechanism is a combination of linear operations and cannot fundamentally increase their non-linear expressive capabilities.
A.2 Approaches with Remarkable Expressiveness but Approximate SO(3)-equivariance
Li et al. [6] proposed a local coordinate strategy to achieve SO(3)-equivariance. This strategy relied on the transformation relationship between local and global coordinates to achieve equivariance, without imposing restrictions on the neural network, allowing the use of network modules that possess non-linear expressiveness. However, this strategy, solely depending on coordinate transformations, only works under holistic rigid rotations of the atomic system. Without exploration of a neural mechanism for equivariance, when the local coordinates experience non-rigid perturbations such as thermal movements or bilayer twists, this strategy offers no guarantees on maintaining symmetry properties; Zitnick et al. [11] decomposed features into weighted sums of spherical harmonic basis functions, each corresponding to a discrete point on a unit sphere, through inner-product operations. Such discrete decomposition enables non-linear operations on the coefficients to enhance expressiveness. In subsequent developments [15, 16, 32], this approach has demonstrated a remarkable capacity to fit complex functions. However, the inner-product operations with discrete basis make such approach degenerate from continuous to discrete rotational equivariance, no longer maintaining strict equivariance to continuous rotational operations, as pointed out by Zhang et al. [1]. Additionally, due to the high complexity of its discrete sampling strategies, tensor orders ($m$) corresponding to large degrees ($l$) are often partially pruned in architectures such as Equiformerv2 [16], further sacrificing the equivariance of features. In summary, although these methods with approximate SO(3)-equivariance achieve high numerical accuracy due to their extensive use of non-linear mechanisms, they are best suited for the applications that primarily require numerical approximations but not for tasks that demand theoretical depth in physics. In some applications, the demands for symmetry are extremely high. Even minor deviations from perfect symmetry can result in incorrect physical results.
Combining these analyses, existing deep learning methods have compromised either strict SO(3)-equivariance or non-linear expressiveness when processing physical systems. In this paper, we develop a novel framework that theoretically combines strict SO(3)-equivariance with the non-linear expressiveness of neural networks. This solution resolves the equivariance-expressiveness dilemma by deeply exploring the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities and their representations.
Appendix B Proof of Theory
B.1 Proof of Theorem 1
Proof.

### v4

@@ 6.1 Experimental Conditions
We use a comprehensive set of metrics to deeply evaluate the performance of deep learning electronic-structure Hamiltonian prediction models. The metrics can be broadly categorized into two major aspects:
• Metrics on accuracy performance. On the databases from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023), we follow Yin et al. (2024) to adopt a set of Mean Absolute Error (MAE) metrics between predicted and ground truth Hamiltonians, including $MAE^{H}_{all}$ for measuring average MAE of all samples and matrix elements, $MAE^{H}_{cha\_s}$ for measuring the MAE of challenging samples where the baseline model performs the worst, $MAE^{H}_{block}$ for measuring the MAE of different basic blocks in the Hamiltonian matrix, and $MAE^{H}_{cha\_b}$ for measuring the MAE on the most challenging Hamiltonian block where the baseline model shows the poorest performance (with the largest $MAE^{H}_{block}$). These metrics comprehensively reflect the accuracy performance, covering not only the average accuracy but also the accuracy on difficult samples and challenging blocks of the Hamiltonian matrices. On the two databases from the QH9 benchmark series, we adopt the metrics introduced by their original paper (Yu et al., 2023a), including MAE of Hamiltonian matrices, which are further subdivided into $MAE^{H}_{all}$ for measuring average MAE, $MAE^{H}_{diag}$ for measuring MAE of Hamiltonian matrix formed by an atom with itself, and $MAE^{H}_{non\_diag}$ for measuring MAE of Hamiltonian matrices formed by different atoms; as well as the MAE ($MAE^{\epsilon}$) of occupied orbital energies $\epsilon$ induced by the predicted Hamiltonians and compared to the ground truth ones, and the cosine similarity ($Sim(\psi)$) between the electronic wavefunctions $\psi$ induced by the predicted and ground truth Hamiltonians. $\epsilon$ and $\psi$ are crucial downstream physical quantities for determining multiple properties of the atomic systems as well as their dynamics, highly reflecting the application values of the Hamiltonian regression model.
• Metrics on acceleration ratios for the convergence of traditional DFT algorithms. Despite the increasing ability of deep learning models to independently handle more electronic-structure computation tasks, there are still applications with extremely high numerical precision requirements and very low tolerance for error, where traditional DFT algorithms must perform the final calculations. In such cases, the predictions from deep models can be used as initial matrices to accelerate the convergence of traditional DFT algorithms. We evaluate the acceleration ratios brought by the proposed method for the convergence of classical DFT algorithms implemented by PySCF (Sun et al., 2018). Specifically, we adopt the two metrics defined in Yu et al. (2023a): the achieved ratio calculates the number of DFT optimization steps taken when initializing with the Hamiltonian matrices predicted by the deep model compared to using initial guess methods like minao and 1e. The error-level ratio measures the number of DFT optimization steps required, starting from random initialization, to reach the same error level as the deep model’s predictions, relative to the total number of steps in the DFT process.
Table 3: The acceleration ratios of QHNet and QHNet+TraceGrad for DFT calculation. Both models are evaluated on a set of $50$ molecules chosen by Yu et al. (2023a), with the mean and standard deviation of the metrics across these samples reported. $\downarrow$ means lower values correspond to better accuracy, while $\uparrow$ means higher values correspond to better performance.
Methods Training DFT Metric Ratio
databases initialization
QHNet (Baseline) QS 1e Achieved ratio $\downarrow$ $0.400\pm 0.030$
Error-level ratio $\uparrow$ $0.620\pm 0.037$
minao Achieved ratio $\downarrow$ $0.715\pm 0.033$
Error-level ratio $\uparrow$ $0.406\pm 0.021$
QD 1e Achieved ratio $\downarrow$ $0.512\pm 0.138$
Error-level ratio $\uparrow$ $0.622\pm 0.048$
minao Achieved ratio $\downarrow$ $0.882\pm 0.217$
Error-level ratio $\uparrow$ $0.406\pm 0.066$
QHNet+TraceGrad QS 1e Achieved ratio $\downarrow$ $\textbf{0.345}\pm 0.038$
Error-level ratio $\uparrow$ $\textbf{0.685}\pm 0.037$
minao Achieved ratio $\downarrow$ $\textbf{0.647}\pm 0.061$
Error-level ratio $\uparrow$ $\textbf{0.466}\pm 0.035$
QD 1e Achieved ratio $\downarrow$ $\textbf{0.440}\pm 0.101$
Error-level ratio $\uparrow$ $\textbf{0.645}\pm 0.046$
minao Achieved ratio $\downarrow$ $\textbf{0.761}\pm 0.167$
Error-level ratio $\uparrow$ $\textbf{0.435}\pm 0.052$
6.2 Results and Analysis
We compare experimental results from two setups: the first one is the experimental results of the baseline SO(3)-equivariant regression model (Gong et al., 2023; Yu et al., 2023b) for Hamiltonian prediction, and the second one is the experimental results of extending the architecture and pipeline of the baseline model through the proposed TraceGrad method, which incorporates non-linear expressiveness into the SO(3)-equivariant features of the baseline model with the gradient operations of SO(3)-invariant non-linear features learned under the supervision of the trace targets. We choose DeepH-E3 (Gong et al., 2023) as the baseline model for databases from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023); and we choose QHNet (Yu et al., 2023b) as the baseline model for databases from the QH9 benchmark series (Yu et al., 2023a). They are the respective state-of-the-art (SOTA) methods with strict SO(3)-equivariance on the corresponding databases.
We list the results of DeepH-E3 and DeepH-E3+TraceGrad in Tables 1 for databases from the DeepH benchmark series, reporting the values of $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$, and $MAE^{H}_{cha\_b}$. The results of DeepH-E3 to be compared are copied from Yin et al. (2024). The results of DeepH-E3+TraceGrad are the average from 10 independent repeated experiments. Regarding the metric of $MAE^{H}_{block}$ for every Hamiltonian block, due to its large data volume, we just present its values for all databases from the DeepH series in Appendix 4. We use the same fixed random seed as adopted by DeepH-E3 for all random processes in experiments on these six databases. As a result, the standard deviation of the Hamiltonian prediction MAE across repeated experiments does not exceed $0.007$ meV for each of the six databases and is negligible.
From results presented in Tables 1, we could find that the proposed TraceGrad method dramatically enhances the accuracy performance of the baseline method DeepH-E3, both on average and for challenging samples and blocks, both on the non-twisted samples and the twisted samples. Specifically, on the corresponding datasets, TraceGrad lowers down the $MAE^{H}_{all}$ and $MAE^{H}_{cha}$ of DeepH-E3 with relative ratios of up to $34\%$ and $35\%$, respectively. Furthermore, from the results included in Appendix 4, TraceGrad significantly improves the performance for the vast majority of basic blocks. Particularly, for the blocks where DeepH-E3 perform the worst, TraceGrad reduces the $MAE$ ($MAE^{H}_{cha\_b}$) by a maximum of $48\%$. The leading performance on the $MG$ and $MM$ databases prepared at three-hundred Kelvin temperature demonstrates the robustness of our method against thermal motion. The high accuracy on the $BB$, $BT$, and $BS$ databases, which have strong SOC effects, indicates our method’s strong capability to model such effects. The excellent performance on the $BG^{t}$, $BB^{t}$, $BT^{t}$, and $BS^{t}$ subsets showcases the method’s superior generalization to twisted structures, which are not present in the training data. The outstanding performance on such samples highlights the good potential for studying twist-related phenomena, a hot research topic that may bring new electrical and transport properties (Cao et al., 2018; Wang et al., 2024a; He et al., 2024). Additionally, the $BG^{t}$, $BB^{t}$, $BT^{t}$, and $BS^{t}$ subsets contain significantly larger unit cells compared to the training set (see Appendix C for statistics of their sizes), yet our method still excels on these subsets as measured by the multiple MAE metrics, demonstrating its good scalability on the sizes of atomic systems it handles.
In Table 2, we present the results of QHNet and QHNet+TraceGrad under the metrics of $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ for the $QS$ and $QD$ databases. The results of QHNet to be compared are taken from their original paper (Yu et al., 2023a), and for the unification of MAE units, we convert the units of MAE from $10^{-6}$ Hartree ($E_{h}$) in the original paper to meV111 $27211.41E_{h}=$ 27211.4 meV. The results of QHNet+TraceGrad are the average from 10 independent repeated experiments. To ensure reproducibility, we use the same fixed random seeds as employed in QHNet for all random processes in the experiments on the $QS$ and $QD$ databases. As a result, the standard deviation of the Hamiltonian prediction MAE across repeated experiments is no greater than $0.009$ meV for both $QS$ and $QD$ and is also negligible.
The results presented in Table 2 demonstrate that the proposed TraceGrad method significantly enhances the accuracy of the baseline QHNet model across all metrics on the $QS$ and $QD$ databases. Specifically, TraceGrad reduces $MAE^{H}_{all}$, $MAE^{H}_{non\_diag}$, $MAE^{H}_{diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ of QHNet with relative reductions of up to 40%, 39%, 40%, 76%, and 17%, respectively, on the corresponding databases. The significant accuracy improvements on the $QS$ database, partitioned using the ’ood’ split strategy (Yu et al., 2023a) without scale overlapping among the training, validation, and testing sets, once again demonstrate the method’s strong generalization capabilities across different scales of atomic systems. Meanwhile, performance on the $QD$ database under the ’mol’ strategy (Yu et al., 2023a), which partitions the training, validation, and testing sets with samples from completely different thermal motion trajectories, highlight our method’s robustness in generalizing to new thermal motion sequences. Furthermore, the substantial improvement in the prediction accuracy of $\epsilon$, i.e., occupied orbital energies crucial for determining electronic properties such as optical characteristics and conductivity in atomic systems, and $\psi$, i.e., the electronic wavefunctions essential for understanding electron distribution and interactions, underscores the potential values of our method for applications like material design, molecular pharmacology, and quantum computing.
We also compare the acceleration ratios brought by the QHNet model and the QHNet+TraceGrad model for the convergence of classical DFT algorithms implemented by PySCF (Sun et al., 2018), with the experimental results reported in Table 3, using the achieved ratio and error-level ratio metrics introduced in Section 6.1. As observed from this table, TraceGrad also brings significant improvements to the baseline model QHNet, notably reducing the achieved ratio and enhancing the error-level ratio, indicating its substantial potential for accelerating traditional DFT algorithms.
Due to the page limits, we present three groups of supplementary experimental results in the Appendices. The first group, included in Appendix H, is the results of ablation study. These fine-grained experiments assess the individual and combined contributions of the two core mechanisms of the proposed TraceGrad method on the six databases from the DeepH series. These mechanisms are the SO(3)-invariant trace supervision mechanism (Trace) at the label level and the gradient-based induction mechanism (Grad) at the representation layer. Results indicate that each mechanism can contribute individually to the performance. Moreover, their combination provides even better performance. The second group of experiments, as included in Appendix I, test the extra time cost caused by TraceGrad compare to the baseline model. As we adopt a light-weight implementation of the non-linear modules introduced in the TraceGrad method, experimental results show that the extra time burden is slight relative to the baseline models DeepH-E3 and QHNet, indicating the good efficiency of our approach. The third group of experiments corresponds to the synergy of our method with an approximately SO(3)-equivariant methodology (Yin et al., 2024). As approximate equivariance is not the main focus of this paper, we just include these experimental results in Appendix J. The experimental results in Appendix J still demonstrate a significant accuracy improvement brought by our approach, further supporting the generality of our methodology.
7 Conclusion
We propose a theoretical and methodological framework to tackle the issue of reconciling non-linear expressiveness with SO(3)-equivariance in deep learning frameworks for physical system modeling, through deeply investigating the mathematical connections between SO(3)-invariant and SO(3)-equivariant quantities, as well as their representations. We first constructs SO(3)-invariant quantities from SO(3)-equivariant regression targets, using them to train informative SO(3)-invariant non-linear representations. From these, SO(3)-equivariant features are derived with gradient operations, achieving non-linear expressiveness while maintaining strict SO(3)-equivariance. We apply our theory and method to the challenging electronic-structure Hamiltonian prediction tasks, achieving dramatic promotions in prediction accuracy across eight benchmark databases. Experimental results demonstrate that this approach not only improves the accuracy of Hamiltonian prediction but also significantly enhances the prediction for downstream physical quantities, and also markedly improves the acceleration ratios for traditional DFT algorithms.
Ethics Statement
This work develops a representation learning method that exhibits strong non-linear expressive capabilities while strictly adhering to SO(3) equivariance. This method has demonstrated superior accuracy in predicting electronic-structure Hamiltonians and related physical quantities, showcasing its potential to accelerate research in materials science and molecular pharmacology. While we recognize that our research area has not yet revealed direct negative social or ethical implications, several issues warrant our vigilance. Currently, although our method yields accurate predictions, the decision-making processes of deep learning systems often lack transparency, hindering a comprehensive understanding of the learning outcomes and limiting our ability to gain deeper insights. We believe it is important to investigate the interpretability of such models, particularly in terms of how they apply physical knowledge in a comprehensible way. Additionally, it is crucial to continually improve the correctness and fairness of deep learning models on this area. Ensuring high-quality and diverse training data, implementing sound model designs, and performing ongoing validation and refinement are necessary to guarantee model accuracy and the broad applicability of their results.
Appendices
A Background
Definition 1.
Group. A set $G$, denoted as $G=\{\dots,g,\dots\}$, equipped with a binary operation denoted as $\cdot$, is called a group if it satisfies the following four conditions:
  1. 1.
    Closure: For any two elements $f,g\in G$, $f\cdot g\in G$.
  2. 2.
    Associativity: For all elements $f,g,h\in G$, the equality $(f\cdot g)\cdot h=f\cdot(g\cdot h)$ holds.
  3. 3.
    Existence of Identity Element: There exists a unique identity element $e\in G$ such that for all $f\in G$, $e\cdot f=f\cdot e=f$.
  4. 4.
    Existence of Inverse Element: For each element $f\in G$, there exists a unique inverse element $f^{-1}\in G$ such that $f\cdot f^{-1}=f^{-1}\cdot f=e$.
Definition 2.
SO(3) Group. The special orthogonal group $\text{SO}(3)$ is the group of all $3\times 3$ orthogonal matrices with determinant 1. Formally, it can be defined as:
$\text{SO}(3)=\{\mathbf{R}\in\mathbb{R}^{3\times 3}\mid\mathbf{R}^{T}\mathbf{R}=I,\,\det(\mathbf{R})=1\}$
where $\mathbf{R}^{T}$ denotes the transpose of $\mathbf{R}$ and $I$ is the $3\times 3$ identity matrix. The elements of $\text{SO}(3)$ represent rotations in three-dimensional Euclidean space.
Definition 3.
Group Representation. A representation of a group $G$ on a tensor space $T(V)$ is a homomorphism $\rho$ from $G$ to the general linear group $GL(T(V))$, the group of all invertible linear transformations on $T(V)$. Here, $T(V)$ represents the tensor space associated with the vector space $V$, encompassing all tensors that can be formed from elements of $V$. The homomorphism $\rho$ can be formalized as:
$\rho:G\to GL(T(V))$
such that for all $g_{1},g_{2}\in G$,
$\rho(g_{1}g_{2})=\rho(g_{1})\rho(g_{2})$
and $\rho(e)=I$, where $e$ is the identity element of $G$, and $I$ is the identity transformation on $T(V)$.
Definition 4.
SO(3) Group Representation. A representation of the special orthogonal group $\text{SO}(3)$ on a tensor space $T(V)$ is a homomorphism $\rho:\text{SO}(3)\to GL(T(V))$, where $GL(T(V))$ denotes the group of all invertible linear transformations on the tensor space $T(V)$. This representation can be expressed using Wigner-D matrices $\mathbf{D}^{l}(\mathbf{R})$, which are defined for a given degree $l$ corresponding to the representation of the rotation matrix $\mathbf{R}\in\text{SO}(3)$:
$\rho(\mathbf{R})=\mathbf{D}^{l}(\mathbf{R}).$
Definition 5.
Equivariance with Respect to a Group. Let $G$ be a group, and let $\rho_{T(V)}:G\to GL(T(V))$ and $\rho_{T(W)}:G\to GL(T(W))$ be representations of $G$ on tensor spaces $T(V)$ and $T(W)$, respectively. A map $f:T(V)\to T(W)$ is said to be equivariant with respect to the group $G$ if the following condition holds:
$f(\rho_{T(V)}(g)\circ v)=\rho_{T(W)}(g)\circ f(v)\quad\text{for all }v\in T(V)\,\,and\,\,g\in G.$
where $\circ$ generally denotes the operation defined on the tensor space.
Definition 6.
Invariance with Respect to a Group. Let $G$ be a group, and let $\rho_{T(V)}:G\to GL(T(V))$ be a representation of $G$ on a tensor space $T(V)$. A function $f:T(V)\to T(W)$ is said to be invariant under the group $G$ if the following condition holds:
$f(\rho_{T(V)}(g)\circ v)=f(v)\quad\text{for all }v\in T(V)\,\,and\,\,g\in G.$
This definition indicates that the function $f$ remains unchanged under the action of the group $G$.
B Proofs of Theorems
Proof of Theorem 1.